A Bibliography of Publications in Biometrics: 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

19 May 2021 Version 1.01

Title word cross-reference

2^{n} [49]. 2 × 2 [92, 111]. 2 × m [51]. 2 × n [108]. 3 × 4 [1]. k [144]. N [107]. t^{2} [181]. U [165].

650 [5].

absorbent [198]. abstract [192]. Abstracts [97]. accommodate [52]. Activities [91]. Adaptation [171]. additional [52]. additivity [29]. adequacy [121]. adjustment [186]. age [35]. age-distribution [35]. all-or-none [189]. alleles [144]. Allocation [40, 10]. almost [167]. among [105, 142]. analyse [148]. Analyses [64, 71]. Analysis [101, 112, 199, 1, 172, 83, 188, 15, 115, 122, 6, 14, 175, 69, 29, 71, 102, 13, 119, 197, 180, 28, 138, 179, 167, 158, 120, 63, 26, 17, 148, 137, 43, 27, 66, 8]. analyzing [164]. Announcements [99]. ANOVA [157, 191]. Application [198, 20, 196]. Applications [93, 55, 58]. Approximate [74, 133, 120]. approximation [121]. Approximations [54, 160]. Architect [83]. arising [162]. arithmetic [118]. aspects [179]. assays [153, 24, 172]. associate [1]. associated [104]. Assumption [165]. Assumption-free [165]. asymmetrical [175, 137]. Asymptotic [187, 27]. asymptotically [70]. attachment [113]. Augmented [41]. Augmenting [52]. Australia [91]. available [109]. Aylmer [76, 78, 83].

B [92]. Back [100]. backgrounds [193]. bacteria [113, 126]. balance [39]. Balanced [149, 174, 63, 26, 166, 40]. based [42, 183, 174, 69, 70]. Bayesian [196]. Behrens [74, 88]. Bennett [92]. best [70]. between [104, 154, 141, 71]. between-plot [141]. bias [186]. binomial [44, 12]. bio [153]. bio-assays [153]. Bioassay [82]. Biological [110, 125, 193]. Biologist [84]. Biometric [98, 86]. Biometry [80]. birth [197, 64, 171]. birth-rank [64]. block [15, 153, 26, 62, 166, 149]. blocks [149]. Blood [89]. Book [95, 92, 96, 94, 93]. branching [107, 57, 171]. Bulgren [93].

C [96, 94]. Calculation [5]. carry [145]. carry-over [145]. case [189]. catch [6]. catchability [147]. categorical [199, 106]. censored [70, 119, 67]. certain [151, 121, 109, 55]. chain [185]. chains [198, 183]. change [72]. Characteristics [146]. characters [13]. chi [187, 33, 110]. chi-square [33, 110]. Chiasma [79]. class [162]. classification [150, 23, 71]. clinical [183, 50]. cliniques [96]. closed [50]. coefficient [168, 133, 129, 190]. coefficients [48, 17]. column [140]. Combination [111, 51]. combinations [46]. Combined [26, 55]. Combining [61, 141]. common [26]. Comparison [195, 128]. comparisons [32, 163, 139]. component [148, 180]. Component-estimation [180]. components [150, 104, 178, 10, 142]. composantes [148]. compounds [4]. Computer [93]. concentration [126]. concept [135]. conditional [3]. conditions [35]. Confidence [117, 68, 19, 167, 133]. confounded [175, 137]. Construction [175, 137]. contagious [109]. Contingency [92, 188, 51, 111, 53]. Contributions [82, 86, 89]. Correction [200]. correlation [53]. Correlations [154, 36]. counts [126]. Covariance [180, 104, 194, 1, 17]. covariances [157, 191]. covariate [138]. Critical [21]. cross [105, 31]. curve [6, 25, 184]. Cyclic [166].

D [92, 96]. dans [95]. data [23, 185, 112, 199, 5, 160, 64, 171, 109, 164, 106, 22]. death [197, 171]. degrees [48, 17]. depend [197]. dependent [119]. Derivation [104]. Design [85, 181, 48, 59, 189, 1]. Designs [145, 15, 183, 174, 153, 175, 41, 72, 49, 26, 62, 140, 166, 149, 173, 137, 50, 40]. detection [16]. Deviations [81]. diagnostic [34]. diallel [31]. differential [176]. diffusion [121]. dilution [16, 24, 126]. dimensional [62]. directional [176]. discrete [151, 160]. discriminant [143, 194, 167]. discriminatoire [148]. discriminatory [148]. disproportionate [71, 112, 22]. distances [143]. distributed [118]. distribution [3, 9, 13, 168, 19, 44, 159, 156, 54, 132, 55, 45, 169, 12, 139, 124, 35, 67]. distributions [116, 151, 121, 142, 133, 109, 114, 125]. dosage [172]. double [105, 72, 47, 60, 149]. doubly [70]. Duncan [21].

effect [194]. effectiveness [186]. effects [145, 178, 39]. efficiencies [42]. efficiency [127, 70]. elementary [40]. elimination [41, 63]. Empirical [36, 127]. enumerations [162]. equal [147]. equations [11]. Equilibria [144]. equilibrium [185, 161]. Equivalence [158]. error [10]. errors [172]. Essais [96]. essentials [115, 122]. estimate [146]. Estimates [136, 150, 42, 70, 36]. Estimating [3, 185, 178, 37, 171, 156, 118]. Estimation [170, 7, 190, 11, 104, 123, 134, 180, 18, 44, 47, 177, 55, 45, 126, 195, 53, 67]. estimator [24]. estimators [60, 165, 61, 65]. evaluation [56]. exact [139]. examples [110]. Expectation [81]. Expectations [157, 182, 191]. Expected [176, 110]. Experiences [84]. experiment [123]. experimental [181, 40]. Experiments [85, 103, 4, 14, 10, 28, 58]. exponential [134, 170, 60, 171, 67]. exponentials [46]. expository [11]. exposure [146]. extension [9]. extreme [156]. extreme-value [156].

Factor [115, 122]. factorial [52, 48, 175, 137]. factorials [15]. factors [52, 197]. failure [146]. family [5]. Fiducial [88]. finite [176, 198, 185]. finite-state [185]. Finney [92]. first [107]. Fisher [76, 82, 116, 74, 91, 86, 194, 78, 90, 87, 89, 83, 84, 85]. Fitting [152, 109, 12, 46, 25]. fixed [178]. Flamant [96]. Foreword [77]. Formation [79]. Four [174]. Four-way [174]. Frank [155]. free [165]. freedom [48, 17]. French [148]. frequencies [110, 190, 22]. frequency [20, 114]. Front [75]. function [194]. functions [182, 143, 116, 131].

G [93]. gamma [133]. general [162]. generalised [156, 132]. Generalization [131, 25, 129]. Generalized [27, 66, 143, 142, 38, 55]. generation [166]. generations [107]. Genetic [79, 161, 36]. Genetics [87, 162, 121, 180, 148]. génétique [148]. graduation [20]. grouped [58]. grouping [149]. Groups [89]. Growth [143, 197, 30, 184, 35]. Growth-invariant [143].

H [94]. heritability [42]. heterogeneity [41]. highly [2]. histocompatability [193]. homogeneity [102, 108, 55, 17]. host [16, 38]. Hsu [92]. Humaines [95]. Human [89]. hybrids [105]. hyperbola [152, 18]. Hypergeometric [66]. hypotheses [37].

IBM [5]. IBM-650 [5]. II [122]. immediate [39]. inbreeding [5, 129, 190]. Incomplete [153, 173, 15, 26, 62, 166, 149]. incompletely [101, 160]. independent [68]. index [56]. indirect [130]. Individual [17, 48, 63]. Inference [160, 101]. information [141]. insects [35]. inter [15]. inter-block [15]. interaction [112, 128, 62, 8]. interpretation [4]. Interrelations [142]. interval [68, 19]. intervals [167]. Intra [15]. Intra- [15]. introduction [115, 122]. invariant [143]. investigation [127]. involving [101, 160]. iterated [67]. iterative [178].

J [92, 96]. jackknife [165].

Kärber [171]. kinetic [66]. kinetics [66]. knowledge [47]. Kodlin [200].

lag [8]. Latin [39, 63]. Latscha [92]. lattice [1]. layout [120]. Lee [116]. Lellouch [96]. level [52]. levels [74]. likelihood [60, 158, 109, 53, 190]. limits [117, 133]. line [7]. linear [187, 46, 194, 199, 164, 27]. logarithmic [45]. logistic [37, 25]. logit [132]. logit-normal [132]. lognormally [118]. lognormally-distributed [118].

M [95, 92]. machine [5]. MANOVA [43]. Marginal [22]. Marine [94]. Markov [198, 183, 185]. Mason [93]. Mass [130]. matched [127, 189]. matching [168]. mating [198]. matrices [194]. Matter [100, 75]. Maximum [57, 60, 107, 158, 109, 53, 190]. mean [157, 191]. means [102, 47, 177, 118]. measurements [14, 154]. Medicine [82, 93]. Memoriam [76]. Memories [90]. method [46, 37, 167, 165, 171, 158, 109, 120, 63, 190]. Méthodes [95]. methods [44, 164, 45]. Microbiology [94]. misclassification [167]. mixed [151, 178, 179]. model [48, 134, 170, 178, 13, 37, 179, 38, 125, 34]. Models [4, 101, 115, 193, 199, 160, 128, 62]. moments [158]. mortality [136, 123]. mosaics [135]. multinomial [73]. Multiple [141, 138, 21, 163]. Multivariate [83, 43, 23, 2, 69]. multiway [22].

nested [71]. News [99]. Neyman [54]. non [103, 73, 164, 27]. non-linear [164, 27]. non-normality [103]. non-parametric [73]. none [189]. nonlinear [131]. normal [176, 116, 70, 68, 19, 160, 177, 132]. normality [103]. note [11, 151, 30]. Notes [89]. numbers [71, 180].

observational [186]. observations [16]. obtaining [167]. Occurrences [146]. offspring [42]. one [41, 17]. one-way [41, 17]. Oppenheimer [94]. Optimum [10, 184]. order [39]. ordered [106]. organism [30]. organisms [124]. other [131, 39]. outlook [196].

P [95, 92]. p.b.i.b [1]. Paired [128, 139]. pairs [189, 163]. parameter [3, 70, 171]. parameters [37, 60, 67]. parametric [73]. parent [42]. part [33]. partial [31]. path [27, 8]. pathogen [38]. patterns [30, 135]. Pearson [116]. Pèpe [95]. percentages [22]. percentiles [156]. period [181]. Perrier [95]. person [146]. Personal [90]. phage [113]. phenotype [190]. plans [52]. plot [141, 28, 173]. Poisson [182, 3, 9, 70, 81, 44, 169, 12]. population [107, 57, 136, 197, 195, 35]. populations [198, 68, 177, 118]. positive [176]. power [187, 55]. practical [54]. precision [65]. preliminary [101]. present [112]. principal [69, 148]. principales [148]. priori [47]. probabilistic [193]. probabilities [185]. Probability [88, 167, 139]. probit [158]. problems [23, 193, 54]. procedure [178, 102, 106]. process [107, 57]. processes [171]. processing [5]. Professor [78]. proportions [187]. Publications [155]. purpose [115, 192].

qualitative [23, 13]. quantal [171, 164, 172, 22]. Quantitative [79, 188, 180]. quarter [49].

R [92, 96]. R. [82, 90, 87, 84]. randomness [73, 135]. range [21, 169]. rank [117, 64, 58]. Rapid [33, 169]. rapports [148]. rates [197, 146]. ratio [117, 68, 146]. ratios [33]. reciprocal [8]. Recognisable [88]. Recombination [79]. records [136]. recruitment [123]. rectangular [152, 18, 1]. reference [7]. regions [7]. Regression [119, 20, 42, 10, 7, 141, 60, 164, 17, 27, 66, 65]. relations [105]. relationship [165, 171, 148]. Relationships [71]. Relative [42]. relatives [104]. removal [136]. removing [186]. reparameterization [131]. repeated [14]. repeated-measurements [14]. replicates [49]. replications [149]. research [122]. residual [39]. Response [48, 188, 159, 164, 172, 125]. responses [189]. results [111]. Review [95, 92, 96, 94, 93]. Robinson [166]. Robustness [103, 108]. role [122]. Ronald [86, 78, 83, 85, 76]. Rouquette [96]. row [140]. rules [195].

s [101]. same [185]. Sample [68, 19, 177, 139, 58]. samples [127, 2, 70, 47, 67]. sampling [36]. scaling [106]. Schwartz [96]. Sciences [95]. scoring [190]. screening [58]. second [59]. Selection [184, 176, 198, 5, 161, 130, 144, 56]. sensitivity [103, 164]. Sensory [32]. separation [2]. sequence [73]. sequences [145]. sequential [183, 195, 50, 58]. serial [24]. Series [81, 16, 175, 55, 45]. ses [148]. set [59]. Sets [88, 176, 154, 102, 63, 140]. several [71, 197, 111]. sexlinked [104]. sib [198]. sibships [13]. sigmoid [131]. sign [163]. signed [117]. Significance [81, 92, 101, 74, 2, 33]. similar [103, 154]. simple [134, 170, 30, 63]. simultaneous [11]. single [16, 123, 105]. single-tagging [123]. singly [116, 70, 67]. Sir [86, 83, 85]. situations [178]. size [107, 57, 68, 19, 177, 195]. slope [7]. small [2, 110]. Society [98]. Some [23, 44, 179, 90, 45, 140, 89, 50, 84, 182, 175, 131, 160, 47, 109, 30, 26, 137, 40, 65]. Son [148]. spaces [139]. Spearman [134, 24]. special [7]. specified [101, 68, 19, 160]. split [28, 173]. split-plot [28]. square [187, 59, 33, 110]. squares [174, 157, 63, 191, 39]. state [185]. statistical [193, 18, 179]. Statistics [86, 84, 34, 165]. Statistiques [95]. Stepwise [69]. Stochastic [113, 125]. structural [11]. studies [186, 5]. Sub [88]. Sub-Sets [88]. subclass [180]. subclassification [186]. subpopulations [176]. Survey [193, 136]. survey-removal [136]. Symposium [94]. synthesi [191]. synthesis [157]. system [38].

Table [92]. Tables [116, 188, 51, 108, 111, 53, 22, 92]. tagging [123]. tail [7]. Test [88, 101, 73, 117, 74, 147, 2, 33, 21, 110, 169, 163]. Testing [92, 193, 102, 37, 32, 17]. tests [103, 187, 108, 128, 58]. their [142]. Theoretical [105, 30]. Theory [79]. thérapeutiques [96]. Three [49, 62, 150, 105, 71, 33]. Three-dimensional [62]. three-part [33]. three-way [150, 105, 71]. Tied [72]. Tied-double-change-over [72]. time [159]. times [125]. Tisserand [95]. Tisserand-Perrier [95]. tracer [4]. transformation [192]. transition [185]. transmission [104]. treatment [8]. treatments [145, 59, 174, 26, 140, 163]. trends [187, 63]. trials [73, 183, 50]. triple [32, 1]. true [192]. Truncated [183, 116, 9, 70]. Two [58, 52, 181, 10, 2, 71, 68, 120, 140]. two-level [52]. two-period [181]. two-sample [58]. two-way [71, 120]. Type [54].

unbiased [167, 61, 65]. underlying [115]. unequal [194, 180]. uniform [35]. units [181, 40]. Unusual [114]. upon [197]. Use [47, 101, 156, 164, 148]. used [146]. using [117, 4, 165, 169]. utilisation [148].

value [156, 130]. values [21]. variability [16]. variable [119]. variables [176, 69]. variance [150, 104, 178, 69, 29, 71, 102, 194, 112, 19, 43]. variance-covariance [194]. Variances [150, 68, 157]. variates [182]. variation [133]. variations [191]. viruses [126]. volume [7].

W [93]. way [150, 174, 105, 41, 71, 120, 17]. width [68, 19]. Wilcoxon [117, 155]. within [176, 141]. within- [141]. without [8]. Work [87].

year [146]. Youden [59, 174].

References

Pasternack:1960:ACT

[1] Bernard S. Pasternack. Analysis of covariance for a 3×4 triple rectangular lattice design (3 associate p.b.i.b.). Biometrics, 16(1):7–18, March 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dempster:1960:STS

[2] A. P. Dempster. A significance test for the separation of two highly multivariate small samples. Biometrics, 16(??):41–50, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cohen:1960:EPC

[3] A. Clifford Cohen, Jr. Estimating the parameter in a conditional Poisson distribution. Biometrics, 16(??):203–211, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cornfield:1960:MIE

[4] Jerome Cornfield, Jesse Steinfeld, and Samuel W. Greenhouse. Models for the interpretation of experiments using tracer compounds. Biometrics, 16 (??):212–234, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hoen:1960:CIF

[5] K. Hoen and A. H. E. Grandage. Calculation on inbreeding in family selection studies on the IBM-650 data processing machine. Biometrics, 16(??):292–296, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chapman:1960:ACC

[6] D. G. Chapman and D. S. Robson. The analysis of a catch curve. Biometrics, 16(??):354–368, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dudzinski:1960:ERS

[7] M. L. Dudzinski. Estimation of regression slope from tail regions with special reference to the volume line. Biometrics, 16(??):399–407, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wright:1960:TRI

[8] Sewall Wright. The treatment of reciprocal interaction, with or without lag, in path analysis. Biometrics, 16(??):423–445, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cohen:1960:ETP

[9] A. Clifford Cohen, Jr. An extension of a truncated Poisson distribution. Biometrics, 16(??):446–450, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

DeBaun:1960:OAR

[10] Robert M. De Baun and Victor Chew. Optimum allocation in regression experiments with two components of error. Biometrics, 16(??):451–463, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Basmann:1960:ENE

[11] R. L. Basmann. An expository note on estimation of simultaneous structural equations. Biometrics, 16(??):464–480, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shumway:1960:FPB

[12] Robert Shumway and John Gurland. Fitting the Poisson binomial distribution. Biometrics, 16(??):522–533, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gittelsohn:1960:MAD

[13] A. M. Gittelsohn. A model for the analysis of the distribution of qualitative characters in sibships. Biometrics, 16(??):534–546, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Danford:1960:ARM

[14] M. B. Danford, Harry M. Hughes, and R. C. McNee. On the analysis of repeated-measurements experiments. Biometrics, 16(??):547–565, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bradley:1960:IIB

[15] R. A. Bradley, R. E. Walpole, and C. Y. Kramer. Intra- and inter-block analysis for factorials in incomplete block designs. Biometrics, 16(??): 566–581, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Armitage:1960:DHV

[16] P. Armitage and G. E. Bartsch. The detection of host variability in a dilution series with single observations. Biometrics, 16(??):582–592, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Robson:1960:IDF

[17] D. S. Robson and G. F. Atkinson. Individual degrees of freedom for testing homogeneity of regression coefficients in a one-way analysis of covariance. Biometrics, 16(??):593–605, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hey:1960:SER

[18] E. N. Hey and M. H. Hey. The statistical estimation of a rectangular hyperbola. Biometrics, 16(??):606–617, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Graybill:1960:SSS

[19] Franklin A. Graybill and Robert D. Morrison. Sample size for a specified width confidence interval on the variance of a normal distribution. Biometrics, 16(??):636–641, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Beuhler:1960:ARF

[20] Robert J. Beuhler. An application of regression to frequency graduation. Biometrics, 16(??):659–670, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Harter:1960:CVD

[21] H. Leon Harter. Critical values for Duncan’s new multiple range test. Biometrics, 16(??):671–685, 1960. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yates:1961:MPM

[22] F. Yates. Marginal percentages in multiway tables of quantal data with disproportionate frequencies. Biometrics, 17(??):1–9, 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cochran:1961:SCP

[23] William G. Cochran and Carl E. Hopkins. Some classification problems with multivariate qualitative data. Biometrics, 17(??):10–32, 1961. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Johnson:1961:SES

[24] Eugene A. Johnson and Byron Wm. Brown, Jr. The Spearman estimator for serial dilution assays. Biometrics, 17(??):79–88, 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nelder:1961:FGL

[25] J. A. Nelder. The fitting of a generalization of the logistic curve. Biometrics, 17(??):89–110, 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pavate:1961:CAB

[26] M. V. Pavate. Combined analysis of balanced incomplete block designs with some common treatments. Biometrics, 17(??):111–119, 1961. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Turner:1961:GAR

[27] Malcolm E. Turner, Robert J. Monroe, and Henry L. Lucas, Jr. Generalized asymptotic regression and non-linear path analysis. Biometrics, 17 (??):120–143, 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Harter:1961:ASP

[28] H. Leon Harter. On the analysis of split-plot experiments. Biometrics, 17(??):144–149, 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Elston:1961:AAV

[29] Robert C. Elston. On additivity in the analysis of variance. Biometrics, 17(??):209–219, 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nelder:1961:NSG

[30] J. A. Nelder. A note on some growth patterns in a simple theoretical organism. Biometrics, 17(??):220–228, 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kempthorne:1961:PDC

[31] O. Kempthorne and R. N. Curnow. The partial diallel cross. Biometrics, 17(??):229–250, 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Park:1961:STT

[32] G. T. Park. Sensory testing by triple comparisons. Biometrics, 17(??): 251–260, 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gates:1961:RCS

[33] Charles E. Gates and Benjamin H. Beard. Rapid chi-square test of significance for three-part ratios. Biometrics, 17(??):261–266, 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

vanWoerkom:1961:SDM

[34] Adrianus J. van Woerkom and Keeve Brodman. Statistics for a diagnostic model. Biometrics, 17(??):299–318, 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Williams:1961:GAD

[35] E. J. Williams. The growth and age-distribution of a population of insects under uniform conditions. Biometrics, 17(??):349–358, 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

VanVleck:1961:ESE

[36] L. D. VanVleck and C. R. Henderson. Empirical sampling estimates of genetic correlations. Biometrics, 17(??):359–371, 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Grizzle:1961:NMT

[37] James E. Grizzle. A new method of testing hypotheses and estimating parameters for the logistic model. Biometrics, 17(??):372–385, 1961. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mode:1961:GMH

[38] C. J. Mode. A generalized model of a host–pathogen system. Biometrics, 17(??):386–404, 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sheehe:1961:LSB

[39] Paul R. Sheehe and Irwin D. J. Bross. Latin squares to balance immediate residual, and other order, effects. Biometrics, 17(??):405–414, 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Williams:1961:AEU

[40] J. S. Williams. Allocation of experimental units in some elementary balanced designs. Biometrics, 17(??):438–446, 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Federer:1961:ADO

[41] Walter T. Federer. Augmented designs with one-way elimination of heterogeneity. Biometrics, 17(??):447–473, 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bohren:1961:REH

[42] B. B. Bohren, H. E. McKean, and Yukio Yamada. Relative efficiencies of heritability estimates based on regression of offspring on parent. Biometrics, 17(??):481–491, 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Smith:1962:MAV

[43] H. Smith, R. Gnanadesikan, and J. B. Hughes. Multivariate analysis of variance (MANOVA). Biometrics, 18(??):22–41, 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Katti:1962:SME

[44] S. K. Katti and John Gurland. Some methods of estimation for the Poisson binomial distribution. Biometrics, 18(??):42–51, 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Patil:1962:SME

[45] G. P. Patil. Some methods of estimation for the logarithmic series distribution. Biometrics, 18(??):68–75, 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cornell:1962:MFL

[46] Richard G. Cornell. A method for fitting linear combinations of exponentials. Biometrics, 18(??):104–113, 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Katti:1962:USP

[47] S. K. Katti. Use of some a priori knowledge in the estimation of means from double samples. Biometrics, 18(??):139–147, 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bose:1962:RMC

[48] R. C. Bose and R. L. Carter. Response model coefficients and the individual degrees of freedom of a factorial design. Biometrics, 18(??):160–171, 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

John:1962:TQR

[49] Peter W. M. John. Three quarter replicates of 2^{n} designs. Biometrics, 18(??):172–184, 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Spicer:1962:SNC

[50] C. C. Spicer. Some new closed sequential designs for clinical trials. Biometrics, 18(??):203–211, 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kincaid:1962:CCT

[51] W. M. Kincaid. The combination of 2×m contingency tables. Biometrics, 18(??):224–228, 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Addelman:1962:AFP

[52] Sidney Addelman. Augmenting factorial plans to accommodate additional two-level factors. Biometrics, 18(??):308–322, 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tallis:1962:MLE

[53] G. M. Tallis. The maximum likelihood estimation of correlation from contingency tables. Biometrics, 18(??):342–353, 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Martin:1962:ANT

[54] Donald C. Martin and S. K. Katti. Approximations to the Neyman Type A distribution for practical problems. Biometrics, 18(??):354–364, 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Patil:1962:HCE

[55] G. P. Patil. On homogeneity and combined estimation for the generalized power series distribution and certain applications. Biometrics, 18(??): 365–374, 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Williams:1962:ESI

[56] J. S. Williams. The evaluation of a selection index. Biometrics, 18(??): 375–393, 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bishir:1962:MPS

[57] John Bishir. Maximum population size in a branching process. Biometrics, 18(??):394–403, 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wilcoxon:1963:TST

[58] Frank Wilcoxon, L. J. Rhodes, and Ralph A. Bradley. Two sequential twosample grouped rank tests with applications to screening experiments. Biometrics, 19(??):58–84, 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Clarke:1963:SST

[59] G. M. Clarke. A second set of treatments in a Youden square design. Biometrics, 19(??):98–104, 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lipton:1963:MLE

[60] S. Lipton and C. McGilchrist. Maximum likelihood estimators of parameters in double exponential regression. Biometrics, 19(??):144–151, 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Seshadri:1963:CUE

[61] V. Seshadri. Combining unbiased estimators. Biometrics, 19(??):163–170, 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Potthoff:1963:TDI

[62] Richard F. Potthoff. Three-dimensional incomplete block designs for interaction models. Biometrics, 19(??):229–263, 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ogilvie:1963:SME

[63] John C. Ogilvie. A simple method for the elimination of individual trends in the analysis of balanced sets of Latin squares. Biometrics, 19(??): 264–272, 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mantel:1963:ABR

[64] Nathan Mantel and Max Halperin. Analyses of birth-rank data. Biometrics, 19(??):324–340, 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Williams:1963:PSU

[65] W. H. Williams. The precision of some unbiased regression estimators. Biometrics, 19(??):352–361, 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Turner:1963:GKR

[66] Malcolm E. Turner, Robert J. Monroe, and Louis D. Homer. Generalized kinetic regression analysis: Hypergeometric kinetics. Biometrics, 19(??): 406–428, 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Winer:1963:EPI

[67] P. Winer. The estimation of the parameters of the iterated exponential distribution from singly censored samples. Biometrics, 19(??):460–464, 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goldman:1963:SSS

[68] Aaron Goldman. Sample size for a specified width confidence interval on the ratio of variances from two independent normal populations. Biometrics, 19(??):465–477, 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dempster:1963:SMA

[69] A. P. Dempster. Stepwise multivariate analysis of variance based on principal variables. Biometrics, 19(??):478–490, 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Doss:1963:EBA

[70] S. A. D. C. Doss. On the efficiency of best asymptotically normal estimates of the Poisson parameter based on singly and doubly truncated or censored samples. Biometrics, 19(??):588–594, 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Federer:1963:RBT

[71] W. T. Federer. Relationships between a three-way classification disproportionate numbers analysis of variance and several two-way classification and nested analyses. Biometrics, 19(??):629–637, 1963. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Federer:1964:TDC

[72] W. T. Federer and G. F. Atkinson. Tied-double-change-over designs. Biometrics, 20(??):168–181, 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bennett:1964:NPT

[73] B. M. Bennett. A non-parametric test for randomness in a sequence of multinomial trials. Biometrics, 20(??):182–190, 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cochran:1964:ASL

[74] William G. Cochran. Approximate significance levels of the Behrens– Fisher test. Biometrics, 20(??):191–195, 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:1964:FM

[75] Anonymous. Front matter. Biometrics, 20(2):??, June 1964. URL https:

//www.jstor.org/stable/2528389.

Anonymous:1964:MRA

[76] Anonymous. In memoriam: Ronald Aylmer Fisher, 1890–1962. Biometrics, 20(2):i–iii, June 1964. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2528390.

Finney:1964:F

[77] D. J. Finney. Foreword. Biometrics, 20(2):237, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528391.

Mahalanobis:1964:PRA

[78] P. C. Mahalanobis. Professor Ronald Aylmer Fisher. Biometrics, 20(2): 238–252, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/2528392.

Fisher:1964:QTG

[79] R. A. Fisher. A quantitative theory of genetic recombination and chiasma formation. Biometrics, 20(2):253–260, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528393.

Fisher:1964:B

[80] R. A. Fisher. Biometry. Biometrics, 20(2):261–264, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528394.

Fisher:1964:SDE

[81] R. A. Fisher. The significance of deviations from expectation in a Poisson series. Biometrics, 20(2):265–272, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528395.

Bliss:1964:RFC

[82] C. I. Bliss. R. A. Fisher’s contributions to medicine and bioassay. Biometrics, 20(2):273–285, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528396.

Rao:1964:SRA

[83] C. Radhakrishna Rao. Sir Ronald Aylmer Fisher — the architect of multivariate analysis. Biometrics, 20(2):286–300, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528397.

Williams:1964:SEB

[84] C. B. Williams. Some experiences of a biologist with R. A. Fisher and statistics. Biometrics, 20(2):301–306, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528398.

Yates:1964:SRF

[85] F. Yates. Sir Ronald Fisher and the design of experiments. Biometrics, 20(2):307–321, June 1964. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2528399.

Finney:1964:SRF

[86] D. J. Finney. Sir Ronald Fisher’s contributions to biometric statistics. Biometrics, 20(2):322–329, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528400.

Mather:1964:RFW

[87] K. Mather. R. A. Fisher’s work in genetics. Biometrics, 20(2):330–342, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528401.

Yates:1964:FPR

[88] F. Yates. Fiducial probability, recognisable sub-sets and Behrens’ test. Biometrics, 20(2):343–360, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528402.

Race:1964:SNF

[89] R. R. Race. Some notes on Fisher’s contributions to human blood groups. Biometrics, 20(2):361–367, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528403.

Mahalanobis:1964:SPM

[90] P. C. Mahalanobis. Some personal memories of R. A. Fisher. Biometrics, 20(2):368–371, June 1964. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2528404.

Cornish:1964:FAA

[91] E. A. Cornish. Fisher’s activities in Australia, 1959–62. Biometrics, 20 (2):372–373, June 1964. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/2528405.

Mountford:1964:BRT

[92] M. D. Mountford. Book review: Tables for Testing Significance in a 2 × 2 Contingency Table. by D. J. Finney, R. Latscha, B. M. Bennett, P. Hsu. Biometrics, 20(2):374, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

2528406.

Searle:1964:BRC

[93] S. R. Searle. Book review: Computer Applications in Medicine. by E. E. Mason, W. G. Bulgren. Biometrics, 20(2):374–375, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528407.

Regier:1964:BRS

[94] H. A. Regier. Book review: Symposium on Marine Microbiology. by C. H. Oppenheimer. Biometrics, 20(2):375, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528408.

Healy:1964:BRM

[95] M. J. R. Healy. Book review: Méthodes Statistiques dans les Sciences Humaines. by P. Pèpe, M. Tisserand-Perrier. Biometrics, 20(2):375–376, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528409.

Nelder:1964:BRE

[96] J. A. Nelder. Book review: Les Essais thérapeutiques cliniques. by D. Schwartz, R. Flamant, J. Lellouch, C. Rouquette. Biometrics, 20(2):376, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528410.

Anonymous:1964:A

[97] Anonymous. Abstracts. Biometrics, 20(2):377–402, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528411.

Anonymous:1964:BS

[98] Anonymous. The Biometric Society. Biometrics, 20(2):403–421, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528412.

Anonymous:1964:NA

[99] Anonymous. News and announcements. Biometrics, 20(2):422–425, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528413.

Anonymous:1964:BM

[100] Anonymous. Back matter. Biometrics, 20(2):426–xii, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/2528414.

Bancroft:1964:AII

[101] T. A. Bancroft. Analysis and inference for incompletely specified models involving the use of preliminary test(s) of significance. Biometrics, 20 (??):427–442, 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gabriel:1964:PTH

[102] K. R. Gabriel. A procedure for testing the homogeneity of all sets of means in analysis of variance. Biometrics, 20(??):459–477, 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Atiqullah:1964:RNN

[103] M. Atiqullah. Robustness to non-normality of tests for sensitivity in similar experiments. Biometrics, 20(??):478–481, 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bohidar:1964:DEV

[104] Neeti R. Bohidar. Derivation and estimation of variance and covariance components associated with covariance between relatives under sexlinked transmission. Biometrics, 20(??):505–521, 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Eberhart:1964:TRA

[105] S. A. Eberhart. Theoretical relations among single, three-way, and double cross hybrids. Biometrics, 20(??):522–539, 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Snell:1964:SPO

[106] E. J. Snell. A scaling procedure for ordered categorical data. Biometrics, 20(??):592–607, 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Adke:1964:MPS

[107] S. R. Adke. The maximum population size in the first N generations of a branching process. Biometrics, 20(??):649–651, 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lewontin:1965:RHT

[108] R. C. Lewontin and J. Felsenstein. The robustness of homogeneity tests in 2 × n tables. Biometrics, 21(??):19–33, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Martin:1965:FCC

[109] D. C. Martin and S. K. Katti. Fitting of certain contagious distributions to some available data by the maximum likelihood method. Biometrics, 21 (??):34–48, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Katti:1965:BES

[110] S. K. Katti and A. N. Sastry. Biological examples of small expected frequencies and the chi-square test. Biometrics, 21(??):49–54, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Radhakrishna:1965:CRS

[111] S. Radhakrishna. Combination of results from several 2 × 2 contingency tables. Biometrics, 21(??):86–98, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gosslee:1965:AVD

[112] David G. Gosslee and Henry L. Lucas. Analysis of variance of disproportionate data when interaction is present. Biometrics, 21(??):115–133, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gani:1965:SPA

[113] J. Gani. Stochastic phage attachment to bacteria. Biometrics, 21(??): 134–139, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Newell:1965:UFD

[114] D. J. Newell. Unusual frequency distributions. Biometrics, 21(??):159– 168, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cattell:1965:FAIa

[115] Raymond B. Cattell. Factor analysis: an introduction to essentials. I. The purpose and underlying models. Biometrics, 21(??):190–215, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cicchinelli:1965:TPL

[116] A. L. Cicchinelli. Tables of Pearson–Lee–Fisher functions of singly truncated normal distributions. Biometrics, 21(??):219–226, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bennett:1965:CLR

[117] B. M. Bennett. Confidence limits for a ratio using Wilcoxon’s signed rank test. Biometrics, 21(??):231–234, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Oldham:1965:EAM

[118] P. D. Oldham. On estimating the arithmetic means of lognormallydistributed populations. Biometrics, 21(??):235–239, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Glasser:1965:RAD

[119] Marvin Glasser. Regression analysis with dependent variable censored. Biometrics, 21(??):300–307, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mathai:1965:AMA

[120] A. M. Mathai. An approximate method of analysis for a two-way layout. Biometrics, 21(??):376–385, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ewens:1965:ADA

[121] W. J. Ewens. The adequacy of the diffusion approximation to certain distributions in genetics. Biometrics, 21(??):386–394, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cattell:1965:FAIb

[122] Raymond B. Cattell. Factor analysis: an introduction to essentials. II. The role of factor analysis in research. Biometrics, 21(??):405–435, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chapman:1965:EMR

[123] Douglas G. Chapman. The estimation of mortality and recruitment from a single-tagging experiment. Biometrics, 21(??):529–542, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Watson:1965:DO

[124] Geoffrey S. Watson. The distribution of organisms. Biometrics, 21(??): 543–550, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shortley:1965:SMD

[125] George Shortley. A stochastic model for distributions of biological response times. Biometrics, 21(??):562–582, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Roberts:1965:ECV

[126] E. A. Roberts and G. G. Coote. The estimation of concentration of viruses and bacteria from dilution counts. Biometrics, 21(??):600–615, 1965. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Billewicz:1965:EMS

[127] W. Z. Billewicz. The efficiency of matched samples: an empirical investigation. Biometrics, 21(??):623–644, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ogilvie:1965:PCM

[128] John C. Ogilvie. Paired comparison models with tests for interaction. Biometrics, 21(??):651–664, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shikata:1965:GIC

[129] Morikazu Shikata. A generalization of the inbreeding coefficient. Biometrics, 21(??):665–681, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Searle:1965:VIS

[130] S. R. Searle. The value of indirect selection. I. Mass selection. Biometrics, 21(??):682–707, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Grosenbaugh:1965:GRS

[131] L. R. Grosenbaugh. Generalization and reparameterization of some sigmoid and other nonlinear functions. Biometrics, 21(??):708–714, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mead:1965:GLN

[132] R. Mead. A generalised logit-normal distribution. Biometrics, 21(??): 721–732, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Linhart:1965:ACL

[133] H. Linhart. Approximate confidence limits for the coefficient of variation of gamma distributions. Biometrics, 21(??):733–738, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cornell:1965:SES

[134] Richard G. Cornell. Spearman estimation for a simple exponential model. Biometrics, 21(??):858–864, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pielou:1965:CRP

[135] E. C. Pielou. The concept of randomness in the patterns of mosaics. Biometrics, 21(??):908–920, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chapman:1965:EMP

[136] Douglas G. Chapman and Garth I. Murphy. Estimates of mortality and population from survey-removal records. Biometrics, 21(??):921–935, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sardana:1965:CAS

[137] M. G. Sardana and M. N. Das. On the construction and analysis of some confounded asymmetrical factorial designs. Biometrics, 21(??):948–956, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Katti:1965:MCA

[138] S. K. Katti. Multiple covariate analysis. Biometrics, 21(??):957–974, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Trawinski:1965:EPD

[139] B. J. Trawinski. An exact probability distribution over sample spaces of paired comparisons. Biometrics, 21(??):986–1000, 1965. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Preece:1966:SRC

[140] D. A. Preece. Some row and column designs for two sets of treatments. Biometrics, 22(??):1–25, 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Duncan:1966:MRC

[141] David B. Duncan and Mackenzie Walser. Multiple regression combining within- and between-plot information. Biometrics, 22(??):26–43, 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Katti:1966:IAG

[142] S. K. Katti. Interrelations among generalized distributions and their components. Biometrics, 22(??):44–52, 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Burnaby:1966:GID

[143] T. P. Burnaby. Growth-invariant discriminant functions and generalized distances. Biometrics, 22(??):96–110, 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tallis:1966:EUS

[144] G. M. Tallis. Equilibria under selection for k alleles. Biometrics, 22(??): 121–127, 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Atkinson:1966:DST

[145] G. F. Atkinson. Designs for sequences of treatments with carry-over effects. Biometrics, 22(??):292–309, 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sheps:1966:CRU

[146] Mindel C. Sheps. Characteristics of a ratio used to estimate failure rates: Occurrences per person year of exposure. Biometrics, 22(??):310–321, 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cormack:1966:TEC

[147] R. M. Cormack. A test for equal catchability. Biometrics, 22(??):330–342, 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rouvier:1966:LCP

[148] R. Rouvier. L’analyse en composantes principales: Son utilisation en génétique et ses rapports avec l’analyse discriminatoire. (French) [Principal component analysis: Its use in genetics and its relationship to discriminatory analysis]. Biometrics, 22(??):343–357, 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Robinson:1966:BIB

[149] J. Robinson. Balanced incomplete block designs with double grouping of blocks into replications. Biometrics, 22(??):368–373, 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Blischke:1966:VEV

[150] W. R. Blischke. Variances of estimates of variance components in a threeway classification. Biometrics, 22(??):553–565, 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cohen:1966:NCD

[151] Clifford A. Cohen, Jr. A note on certain discrete mixed distributions. Biometrics, 22(??):566–572, 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bliss:1966:FRH

[152] C. I. Bliss and A. T. James. Fitting the rectangular hyperbola. Biometrics, 22(??):573–602, 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Das:1966:IBD

[153] M. N. Das and G. A. Kulkarni. Incomplete block designs for bio-assays. Biometrics, 22(??):706–729, 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

DeGroot:1966:CBS

[154] M. H. DeGroot and C. C. Li. Correlations between similar sets of measurements. Biometrics, 22(??):781–790, 1966. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Karas:1967:PFW

[155] James Karas and I. Richard Savage. Publications of Frank Wilcoxon (1892–1965). Biometrics, 23(??):1–10. (1 plate), 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Maritz:1967:UGE

[156] J. S. Maritz and A. H. Munro. On the use of the generalised extreme-value distribution in estimating extreme percentiles. Biometrics, 23(??):79–103, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hartley:1967:EVC

[157] H. O. Hartley. Expectations, variances and covariances of ANOVA mean squares by “synthesis”. Biometrics, 23(??):105–114, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mantel:1967:EML

[158] Nathan Mantel and Samuel W. Greenhouse. Equivalence of maximum likelihood and the method of moments in probit analysis. Biometrics, 23 (??):154–157, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kodlin:1967:NRT

[159] D. Kodlin. A new response time distribution. Biometrics, 23(??):227–239, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [200].

Kale:1967:ISI

[160] B. K. Kale and T. A. Bancroft. Inference for some incompletely specified models involving normal approximations to discrete data. Biometrics, 23 (??):335–348, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:1967:GEU

[161] C. C. Li. Genetic equilibrium under selection. Biometrics, 23(??):397–484, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bennett:1967:GCE

[162] J. H. Bennett. A general class of enumerations arising in genetics. Biometrics, 23(??):517–537, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rhyne:1967:MCS

[163] A. L. Rhyne and R. G. D. Steel. A multiple comparisons sign test: All pairs of treatments. Biometrics, 23(??):539–549, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Moore:1967:UNL

[164] R. H. Moore and R. K. Zeigler. The use of non-linear regression methods for analyzing sensitivity and quantal response data. Biometrics, 23(??): 563–566, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mantel:1967:AFE

[165] Nathan Mantel. Assumption-free estimators using U statistics and a relationship to the jackknife method. Biometrics, 23(??):567–571, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Preece:1967:CGR

[166] D. A. Preece. Cyclic generation of Robinson’s balanced incomplete block designs. Biometrics, 23(??):574–578, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lachenbruch:1967:AUM

[167] Peter A. Lachenbruch. An almost unbiased method of obtaining confidence intervals for the probability of misclassification in discriminant analysis. Biometrics, 23(??):639–645, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goodall:1967:DMC

[168] D. W. Goodall. The distribution of the matching coefficient. Biometrics, 23(??):647–656, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pettigrew:1967:RTP

[169] Hugh M. Pettigrew and William C. Mohler. A rapid test for the Poisson distribution using the range. Biometrics, 23(??):685–692, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cornell:1967:ESE

[170] Richard G. Cornell and Janace A. Speckman. Estimation for a simple exponential model. Biometrics, 23(??):717–737, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mantel:1967:AKM

[171] Nathan Mantel. Adaptation of Kärber’s method for estimating the exponential parameter from quantal data, and its relationship to birth, death, and branching processes. Biometrics, 23(??):739–746, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Patwary:1967:AQR

[172] Kamini M. Patwary and K. David C. Haley. Analysis of quantal response assays with dosage errors. Biometrics, 23(??):747–760, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Robinson:1967:ISP

[173] J. Robinson. Incomplete split plot designs. Biometrics, 23(??):793–802, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Clarke:1967:FWB

[174] G. M. Clarke. Four-way balanced designs based on Youden squares with 5, 6, or 7 treatments. Biometrics, 23(??):803–812, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Das:1967:CAS

[175] M. N. Das and P. S. Rao. Construction and analysis of some new series of confounded asymmetrical factorial designs. Biometrics, 23(??):813–822, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Allaire:1967:ESD

[176] F. R. Allaire. Expected selection differential for positive directional selection on normal variables within sets of finite subpopulations. Biometrics, 23(??):842–845, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Likes:1967:SSE

[177] Jiř´ ı Likeš. Sample size for the estimation of means of normal populations. Biometrics, 23(??):846–849, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cunningham:1968:IPE

[178] E. P. Cunningham and C. R. Henderson. An iterative procedure for estimating fixed effects and variance components in mixed model situations. Biometrics, 24(??):13–25, 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Koch:1968:SAS

[179] Gary G. Koch and Pranab Kumar Sen. Some aspects of the statistical analysis of the “mixed model”. Biometrics, 24(??):27–48, 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Grossman:1968:CAU

[180] M. Grossman and G. A. E. Gall. Covariance analysis with unequal subclass numbers: Component-estimation in quantitative genetics. Biometrics, 24 (??):49–59, 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Balaam:1968:TPD

[181] L. N. Balaam. A two-period design with t^{2} experimental units. Biometrics, 24(??):61–73, 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bartko:1968:ESF

[182] John J. Bartko, Samuel W. Greenhouse, and Clifford S. Patlak. On expectations of some functions of Poisson variates. Biometrics, 24(??):97–102, 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Choi:1968:TSD

[183] S. C. Choi. Truncated sequential designs for clinical trials based on Markov chains. Biometrics, 24(??):159–168, 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tallis:1968:SOG

[184] G. M. Tallis. Selection for an optimum growth curve. Biometrics, 24(??): 169–177, 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cohen:1968:EET

[185] Joel E. Cohen. On estimating the equilibrium and transition probabilities of finite-state Markov chain from the same data. Biometrics, 24(??): 185–187, 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cochran:1968:EAS

[186] W. G. Cochran. The effectiveness of adjustment by subclassification in removing bias in observational studies. Biometrics, 24(??):295–313, 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chapman:1968:APC

[187] D. G. Chapman and Jun mo Nam. Asymptotic power of chi square tests for linear trends in proportions. Biometrics, 24(??):315–327, 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bhapkar:1968:ACT

[188] V. P. Bhapkar. On the analysis of contingency tables with a quantitative response. Biometrics, 24(??):329–338, 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Miettinen:1968:MPD

[189] Olli S. Miettinen. The matched pairs design in the case of all-or-none responses. Biometrics, 24(??):339–352, 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yasuda:1968:EIC

[190] Norikazu Yasuda. Estimation of the inbreeding coefficient from phenotype frequencies by a method of maximum likelihood scoring. Biometrics, 24 (??):915–935, 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rao:1968:EVC

[191] J. N. K. Rao. On expectations, variations, and covariances of ANOVA mean squares by ‘synthesis’. Biometrics, 24(??):963–978, 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tukey:1968:TPT

[192] John W. Tukey. The true purpose of transformation (abstract 1545). Biometrics, 24(??):1041, ???? 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Elandt-Johnson:1969:SHT

[193] Regina C. Elandt-Johnson. Survey of histocompatability testing. Biological backgrounds, probabilistic and statistical models and problems. Biometrics, 25(??):207–283, 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gilbert:1969:EUV

[194] Ethel S. Gilbert. The effect of unequal variance-covariance matrices on Fisher’s linear discriminant function. Biometrics, 25(??):505–515, 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Samuel:1969:CSR

[195] Ester Samuel. Comparison of sequential rules for estimation of the size of a population. Biometrics, 25(??):517–527, 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cornfield:1969:BOA

[196] Jerome Cornfield. The Bayesian outlook and its application. Biometrics, 25(??):617–642, 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goodman:1969:APG

[197] Leo A. Goodman. The analysis of population growth when the birth and death rates depend upon several factors. Biometrics, 25(??):659–681, 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bosso:1969:AFA

[198] José A. Bosso, Oscar M. Sorarrain, and Ewald A. Favret. Application of finite absorbent Markov chains to sib mating populations with selection. Biometrics, 25(1):17–26, 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Grizzle:1969:ACD

[199] James E. Grizzle, C. Frank Starmer, and Gary G. Koch. Analysis of categorical data by linear models. Biometrics, 25(3):489–504, 1969. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:CK

[200] Anonymous. Correction of Kodlin (1967). Biometrics, 61(1):311, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [159].